OpenStudy (anonymous):

a sanack company wants to make a mixture of peanuts and rasins to sell for $5 per pound. The have 50 pounds of rasins that cost $2 per pound. Peanuts sell for $7 per pund. how many pounds of peanuts should they add to make a mixture that sells for $5 per pound

OpenStudy (anonymous):

@kropot72 do you know how to solve this

OpenStudy (kropot72):

Let the number of pounds of peanuts to be added = p (50 * 2) + (p * 7) = 5(50 + p) 100 + 7p = 250 + 5p Can you solve the equation to find p?

OpenStudy (anonymous):

yea hang on

OpenStudy (anonymous):

75

OpenStudy (kropot72):

Good work! Your answer is correct.

OpenStudy (anonymous):

cool thanks man

OpenStudy (kropot72):

You're welcome :)